A screw gauge has 100 divisions on circular scale and pitch is 0.5 mm. When we fix a wire between it's jaws, it reads 2 divisions on main scale and 48 divisions on circular scale. If the screw gauge has no zero error, radius of wire is 
(1) 2.48 mm
(2) 1.24 mm
(3) 0.62 mm
(4) 0.31 mm

Grade:12th pass

1 Answers

D Vishwaksen Reddy
3 years ago
Pitch=0.5mm
No. Of divisions = 100
Least count=0.5/100=0.005
M.S.R=2
n=48
d=MSR+n×Least count
d=2+48×0.005=2.24mm